### v2

# Linear Diffusion Networks: Harnessing Diffusion Processes for Global Interactions

Abstract
Diffusion kernels capture global dependencies. We present Linear Diffusion Networks (LDNs), a novel architecture that reinterprets sequential data processing as a unified diffusion process. Our model integrates adaptive diffusion modules with localized nonlinear updates and a diffusion-inspired attention mechanism. This design enables efficient global information propagation while preserving fine-grained temporal details. LDN overcomes the limitations of conventional recurrent and transformer models by allowing full parallelization across time steps and supporting robust multi-scale temporal representations. Experiments on benchmark sequence modeling tasks demonstrate that LDN delivers superior performance and scalability, setting a new standard for global interaction in sequential data.
1 Introduction
Sequence modeling lies at the core of numerous applications, ranging from natural language processing to time-series forecasting. Traditional recurrent neural networks (RNNs) such as LSTM Hochreiter and Schmidhuber (1997) have shown impressive capabilities; however, their sequential nature limits parallelization and hampers long-range dependency modeling. Recent transformer models Vaswani et al. (2017) have addressed some of these issues by leveraging self-attention to capture global interactions, but they often incur high computational costs and struggle with subtle temporal dynamics.
In response to these challenges, we propose LDN, a novel recurrent architecture that models temporal evolution as a diffusion process. By reinterpreting hidden state updates as a blend of gradual diffusion and local nonlinear transformations, our method naturally propagates information across all time steps. This innovative approach not only facilitates full parallelization but also ensures that each token’s representation is enriched by contributions from the entire sequence.
The key contributions of our work are threefold:
  • We formulate a unified diffusion framework that combines continuous, diffusive updates with discrete attention mechanisms, thereby achieving both global interaction and local sensitivity.
  • We provide theoretical guarantees that our diffusion operations preserve global interactions, ensuring that every token contributes to the final representation.
  • We empirically demonstrate the efficacy of LDN on diverse sequential tasks, where it outperforms traditional RNNs and recent transformer-based models in terms of efficiency and performance.
By bridging the gap between efficient computation and robust representation learning, LDN represents a significant step forward in the field of sequential modeling.
2 Related Work
In this section, we review recent advances in sequential modeling with a focus on linear methods. We begin with a discussion of linear recurrent architectures and their variants, then move on to transformer linear methods and other efficient attention mechanisms, including very recent approaches. Finally, we highlight how our diffusion-based method combines the strengths of these approaches while overcoming their limitations.
2.1 Linear RNNs and Their Variants
Classical recurrent neural networks (RNNs) such as Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) and Gated Recurrent Units (GRU) have long been the backbone of sequence modeling. However, their inherently sequential processing limits parallelization and impedes modeling of long-range dependencies. In response, researchers have developed linear RNNs that simplify recurrence dynamics to enable efficient parallel computation and alleviate vanishing gradients. Further improvements have been proposed with architectures such as the Independently Recurrent Neural Network (IndRNN) Li et al. (2018a), which enhances stability in deep recurrent networks. While these methods boost computational efficiency, they often lack the capacity to capture the rich nonlinear dynamics observed in complex sequences.
2.2 Transformer Linear Methods and Efficient Attention
Transformers Vaswani et al. (2017) revolutionized sequential modeling by leveraging self-attention to capture global dependencies. However, the quadratic complexity of full self-attention has spurred the development of more efficient variants. Linear attention methods Katharopoulos et al. (2020); Choromanski et al. (2021) approximate the attention mechanism to achieve linear complexity, while approaches such as Linformer Wang et al. (2020) and Reformer Kitaev et al. (2020) use low-rank or reversible mechanisms to further reduce computational overhead. Recent methods like Big Bird Zaheer et al. (2020) and Sparse Transformers Child et al. (2019) introduce sparsity to scale to longer sequences, and the Nyströmformer Xiong et al. (2021) employs a sampling-based approach for efficiency. Furthermore, state-of-the-art innovations such as the Routing Transformer Roy et al. (2021), FlashAttention Dao et al. (2022), and CosFormer Zhou et al. (2022) continue to push the envelope on both performance and computational speed. Although these methods offer impressive scalability, they often face trade-offs between capturing smooth temporal evolution and modeling abrupt dynamics.
2.3 Diffusion-Based Approaches and Our Contributions
An emerging line of work has begun to explore diffusion-based formulations in sequence modeling. Diffusion-Convolutional Neural Networks Atwood and Towsley (2016) and Diffusion-Convolutional Recurrent Neural Networks Li et al. (2018b) have demonstrated strong performance in graph-based and spatiotemporal forecasting tasks. Drawing inspiration from spectral graph theory Chung (1997), our proposed LDN reinterprets temporal evolution as a diffusion process. By integrating gradual diffusion updates with localized nonlinear transformations, our model achieves both global information propagation and fine-grained local dynamics. Unlike conventional linear RNNs or transformer linear methods, our approach naturally blends the strengths of efficient computation and robust representation learning, resulting in a model that can capture multi-scale temporal dependencies more effectively.
Overall, our diffusion-based framework represents a powerful alternative that unifies the benefits of linear dynamics and global attention, positioning it as a promising solution for complex sequential tasks in modern applications.
3 Methods
3.1 Overview
We introduce the Diffusion Recurrent Neural Network (LDN), a novel architecture that reinterprets temporal information sharing as a unified diffusion process. LDN integrates adaptive stability mechanisms, multi-scale updates, and innovative attention modules into a single framework. This design enables full parallelization across time steps while simultaneously capturing both gradual (diffusive) and abrupt (attentional) dependencies, thereby overcoming key limitations of conventional recurrent models.
3.2 Input Encoding and Positional Information
Given an input sequence
$X=\{x_{1},x_{2},\dots,x_{T}\},$
each token $x_{t}$ is first embedded into a $d$-dimensional vector. To incorporate temporal order, a positional encoding is added:
$h_{t}^{(0)}=\text{Embed}(x_{t})+\text{PosEnc}(t),\quad t=1,\dots,T.$
Thus, the initial hidden state matrix is defined as
$H^{(0)}\in\mathbb{R}^{T\times d}.$
3.3 Unified Multi-Component Update: Diffusion, Local, and Linear Attentional Modules
LDN evolves its hidden states over $L$ layers using a unified update rule that integrates three complementary components:
$h_{t}^{(\ell)}=h_{t}^{(\ell-1)}+\Delta h_{t}^{(\ell)},$
with the increment decomposed as
$\Delta h_{t}^{(\ell)}=\underbrace{\delta t\cdot\sum_{s=1}^{T}K_{ts}\Bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\Bigr{)}}_{\text{Diffusion Module}}+\underbrace{F\bigl{(}x_{t},h_{t}^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Linear Attention Module}}.$
Adaptive Time-Step $\delta t$:
A (potentially learnable) parameter that governs the diffusion rate. Adaptive time stepping or implicit integration (e.g., Crank–Nicolson) can be employed to ensure numerical stability.
Learnable Diffusion Kernel $K$:
Parameterized to capture temporal causality and locality:
$K_{ts}=\phi(t-s)\cdot g(|t-s|),$
where $\phi(t-s)$ enforces temporal order via asymmetry and $g(|t-s|)$ (e.g., a Gaussian decay) attenuates the influence of distant time steps. Normalization (or softplus activation) is applied to maintain stability.
Local Update $F$:
A fine-grained update function (such as a gating or residual module) that preserves crucial local information and mitigates over-smoothing.
Linear Attention Module $A_{\text{lin}}$:
In contrast to more elaborate attention schemes, we employ a simple linear attention mechanism. Specifically, we first project the hidden states into query, key, and value spaces:
$Q=H^{(\ell-1)}W_{Q},\quad K=H^{(\ell-1)}W_{K},\quad V=H^{(\ell-1)}W_{V},$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices. Next, after optionally applying an elementwise non-linearity $\phi(\cdot)$ (e.g., ReLU), the attention output is computed as
$A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}=\frac{\phi(Q)\left(\phi(K)^{\top}V\right)}{\phi(Q)\left(\phi(K)^{\top}\mathbf{1}\right)},$
where $\mathbf{1}$ is a vector of ones (ensuring row-wise normalization). This formulation yields a weighted sum over all tokens in a fully linear (and scalable) manner.
In compact matrix form, the layer-wise update for all time steps becomes:
$H^{(\ell)}=H^{(\ell-1)}+\delta t\cdot\Bigl{(}K\,H^{(\ell-1)}-\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\Bigr{)}+F(X,H^{(\ell-1)})+A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}.$
3.4 Multi-Scale Diffusion and Residual Connections
To enhance model expressiveness and prevent over-smoothing, LDN incorporates:
  • Multi-Scale Diffusion: Each layer may adopt distinct diffusion rates $\delta t$ or kernel parameters, enabling the modeling of dynamics at multiple temporal scales.
  • Residual Connections: Skip connections—either between layers or from the input to deeper layers—help preserve high-frequency details and facilitate stable training.
3.5 Parallelization and Temporal Dynamics
The diffusion update in LDN operates simultaneously on the entire hidden state matrix $H^{(\ell-1)}$, enabling full parallelization across time steps. Combined with stacked layers that integrate adaptive diffusion, local updates, and linear attention, the model effectively captures both local and global temporal dependencies.
3.6 Output Decoding
After $L$ layers of processing, the final hidden state matrix $H^{(L)}$ contains enriched representations for each time step. Depending on the downstream task:
  • For sequence-to-sequence tasks, $H^{(L)}$ is passed to a dedicated decoding module to generate the output sequence.
  • For sequence classification, a pooling operation (e.g., mean, max, or attention-based pooling) aggregates the temporal dimension before the result is passed to a classification head.
3.7 Training Procedure
LDN is trained end-to-end using backpropagation with task-specific loss functions (e.g., cross-entropy for classification). Key training details include:
  • Optimization: The Adam optimizer is employed with an appropriate learning rate schedule.
  • Stability Constraints: The diffusion kernel $K$ is regularized (via normalization or softplus activation), and adaptive time stepping or implicit integration methods are applied to ensure a stable diffusion process.
  • Regularization: Standard techniques (e.g., dropout, weight decay) are applied to promote smooth evolution of hidden states and prevent over-diffusion.  [^1] Code available at https://github.com/rubberduck529/LDN/
Figure 1: LDN: A unified architecture that integrates diffusion, local updates, and linear attention for parallel token processing.
4 Theory
In this section, we rigorously establish that the diffusion kernel in LDN inherently captures global dependencies across the entire input sequence. We show that, under mild conditions on the learnable kernel $K$, repeated application of the diffusion update leads to a mixing of information such that every output token is influenced by every input token.
4.1 Global Dependency via Iterated Diffusion
To simplify the discussion, consider the pure diffusion update (ignoring for now the local update $F$ and the linear attention module $A_{\text{lin}}$). At each layer $\ell$, the diffusion module updates the hidden state as
$h_{t}^{(\ell)}=h_{t}^{(\ell-1)}+\delta t\cdot\sum_{s=1}^{T}K_{ts}\Bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\Bigr{)}.$
In matrix form, letting
$H^{(\ell)}=\begin{bmatrix}h_{1}^{(\ell)}\\ h_{2}^{(\ell)}\\ \vdots\\ h_{T}^{(\ell)}\end{bmatrix}\quad\text{and}\quad\mathbf{1}=\begin{bmatrix}1\\ 1\\ \vdots\\ 1\end{bmatrix},$
we can write the update as
$H^{(\ell)}=\left(I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right)\right)H^{(\ell-1)}.$
Define the effective diffusion operator as
$A\triangleq I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}.$
After $L$ layers, the cumulative effect of the diffusion process is
$H^{(L)}=A^{L}H^{(0)}.$
4.2 The Global Dependency Theorem
Theorem 1 (Global Dependency Theorem).
Assume that:
  1. 1.
    The diffusion kernel $K$ satisfies $K_{ts}\geq 0$ for all $t,s$.
  2. 2.
    The directed graph $G(K)$ induced by the nonzero entries of $K$ is strongly connected; that is, for any pair of time indices $i$ and $j$, there exists a sequence $\{i=i_{0},i_{1},\dots,i_{k}=j\}$ with $K_{i_{m+1}i_{m}}>0$ for all $m$. (This is an explicit assumption on the design of $K$; note that while $K$ may contain zero entries, its nonzero pattern must ensure strong connectivity.)
Then, there exists a positive integer $L$ (whose value depends on the structure of $K$) such that the $(i,j)$-th entry of the effective diffusion operator $A^{L}$ is strictly positive for all $i,j$:
$[A^{L}]_{ij}>0.$
Equivalently, the hidden state $h_{i}^{(L)}$ depends on every input $h_{j}^{(0)}$ via
$h_{i}^{(L)}=\sum_{j=1}^{T}[A^{L}]_{ij}\,h_{j}^{(0)},$
and
$\frac{\partial h_{i}^{(L)}}{\partial h_{j}^{(0)}}=[A^{L}]_{ij}>0.$
Thus, the diffusion process captures global dependencies.
Proof.
Starting from the diffusion update,
$H^{(L)}=A^{L}H^{(0)},$
where
$A\triangleq I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right),$
we note that $A$ is a perturbation of the identity matrix by a non-negative matrix. Under the stated assumptions, the induced directed graph is strongly connected, so $A$ is irreducible. By a classical result from the Perron–Frobenius theory for irreducible non-negative matrices, there exists a positive integer $L$ such that $A^{L}$ has strictly positive entries, i.e., $[A^{L}]_{ij}>0$ for all $i,j$.
This shows that every output token $h_{i}^{(L)}$ is influenced by every input token $h_{j}^{(0)}$, as required. ∎
4.3 Implications and Discussion
The theorem formalizes the intuition that even if the diffusion kernel $K$ is designed with local attenuation (for example, via a Gaussian decay $g(|t-s|)$ or directional bias $\phi(t-s)$), the strong connectivity assumption guarantees that there is a nonzero path—possibly through intermediate tokens—between any two tokens in the sequence. However, two important points warrant discussion:
  1. 1.
    Quantitative Bound on $L$: The theorem guarantees the existence of a positive integer $L$ such that $[A^{L}]_{ij}>0$ for all $i,j$, but it does not provide a practical upper bound on $L$. In applications, if the graph induced by $K$ is only sparsely connected, the required $L$ might be very large (potentially on the order of thousands), which could limit the practical utility of the pure diffusion process in isolation.
  2. 2.
    Ensuring Strong Connectivity: Although $K$ is allowed to contain zero entries (as per assumption 1), the theorem’s guarantee relies on the assumption that the pattern of nonzero entries in $K$ induces a strongly connected graph. This is a design or learning constraint: one must ensure that despite local attenuation, there are sufficient nonzero connections such that every token is (indirectly) connected to every other token.
In practice, LDN mitigates these concerns by integrating the diffusion module with both a local update $F$ and a linear attention module $A_{\text{lin}}$. These additional components provide alternative pathways for global information propagation, thereby compensating for scenarios where the required $L$ for the diffusion process alone might be impractically high.
Thus, while the theorem establishes a foundational property of the iterative diffusion process, its practical efficacy in capturing global dependencies within LDN is further enhanced by the complementary design of the overall architecture.
4.4 Stability and Convergence of the Diffusion Process
We now analyze the stability of the pure diffusion update, isolating it from the local and attentional components. Recall that the pure diffusion update is given by
$H^{(\ell)}=\left(I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right)\right)H^{(\ell-1)},$
where $K\in\mathbb{R}^{T\times T}$ is the learnable diffusion kernel with $K_{ts}\geq 0$ for all $t,s$ and $\delta t$ is the adaptive time-step parameter.
Theorem (Stability of the Diffusion Update):
Assume that the kernel $K$ satisfies $K_{ts}\geq 0$ for all $t,s$ and that the time-step $\delta t$ is chosen such that
$\delta t\leq\min_{t\in\{1,\dots,T\}}\frac{1}{\sum_{s=1}^{T}K_{ts}}.$
Then, for each layer $\ell$, the diffusion update is a convex combination of the hidden states at layer $\ell-1$. In particular, for every $t$,
$h_{t}^{(\ell)}=\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)h_{t}^{(\ell-1)}+\delta t\sum_{s=1}^{T}K_{ts}\,h_{s}^{(\ell-1)},$
with coefficients satisfying
$\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)\geq 0\quad\text{and}\quad\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)+\delta t\sum_{s=1}^{T}K_{ts}=1.$
Consequently, the diffusion update is non-expansive with respect to the $\ell_{2}$ norm:
$\|H^{(\ell)}\|_{2}\leq\|H^{(\ell-1)}\|_{2}.$
Repeated application of this update thus ensures that the propagation of hidden states remains stable and does not amplify perturbations.
Proof:
For each time step $t$, rewrite the update as
$h_{t}^{(\ell)}=\alpha_{t}\,h_{t}^{(\ell-1)}+\delta t\sum_{s=1}^{T}K_{ts}\,h_{s}^{(\ell-1)},$
where
$\alpha_{t}\triangleq 1-\delta t\sum_{s=1}^{T}K_{ts}.$
By the choice of $\delta t$, we have $\alpha_{t}\geq 0$ and the sum of the coefficients is
$\alpha_{t}+\delta t\sum_{s=1}^{T}K_{ts}=1.$
Thus, $h_{t}^{(\ell)}$ is a convex combination of $\{h_{s}^{(\ell-1)}\}_{s=1}^{T}$. Because convex combinations are non-expansive with respect to the $\ell_{2}$ norm, it follows that
$\|h_{t}^{(\ell)}\|_{2}\leq\max_{s}\|h_{s}^{(\ell-1)}\|_{2}.$
Aggregating over all time steps yields
$\|H^{(\ell)}\|_{2}\leq\|H^{(\ell-1)}\|_{2}.$
This completes the proof. $\blacksquare$
Implications:
This stability guarantee is crucial for the practical training of LDN. By ensuring that the hidden state norms do not increase with layer depth, the theorem alleviates concerns over exploding gradients and uncontrolled error propagation. In tandem with the global mixing properties established earlier, this result underpins the balanced design of LDN—merging robust stability with comprehensive global information propagation.
5 Experiments
In this section, we evaluate LDN on multiple benchmarks, comparing its performance against the Vision Transformer (ViT) baseline Dosovitskiy et al. (2020). We report results on both the large-scale ImageNet dataset Deng et al. (2009) and the CIFAR-10 dataset cif . In addition, we assess the performance of LDN on the GLUE benchmark Wang et al. (2018) and conduct ablation studies to examine the contributions of LDN’s individual components.
5.1 Datasets and Experimental Setup
For ImageNet, we follow standard preprocessing and augmentation strategies as described in Deng et al. (2009). For CIFAR-10, we adopt typical data augmentation techniques (e.g., random cropping and horizontal flipping) to improve generalization. All models are trained using the Adam optimizer Kingma and Ba (2014) with a cosine annealing learning rate schedule, and experiments were conducted on NVIDIA GPUs under comparable hardware settings.
Pretraining for GLUE.
In order to achieve competitive performance on natural language tasks, LDN is first pretrained on a large-scale unlabeled corpus. Following the setups of BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), we combine English Wikipedia and the BookCorpus to form our pretraining dataset. We employ a masked language modeling task augmented with a novel diffusion-based objective that iteratively refines token representations by propagating contextual information. This approach is designed to capture both local and global dependencies effectively. Pretraining is carried out for 1 million steps with a batch size of 256, resulting in a model of approximately 125 million parameters—comparable to both BERT-base and RoBERTa-base—to ensure a fair comparison during fine-tuning on downstream GLUE tasks.
5.2 Results on ImageNet
Table 1 summarizes the performance comparison between ViT and LDN across different model scales (Base, Large, and Huge). Notice that LDN not only achieves higher Top-1 and Top-5 accuracies but does so with fewer parameters and lower computational cost (FLOPs).
Table 1: Comparison of ViT and LDN on ImageNet.
Variant Architecture Params (M) FLOPs (GMac) Top-1 (%) Top-5 (%)
Base ViT 86 17.6 82.5 96.0
LDN 52 10.6 83.1 96.6
Large ViT 307 61.6 84.2 97.0
LDN 181 36.7 84.8 97.3
Huge ViT 632 120.5 84.8 97.3
LDN 373 75.4 85.3 97.8
5.3 Results on CIFAR-10
We further evaluate the models on CIFAR-10. As shown in Table 2, LDN achieves a test accuracy of 80.9%, outperforming the ViT baseline which attains 77.5%. Figure 2 plots the training accuracy over epochs, illustrating that LDN converges faster and reaches a higher accuracy than ViT.
Table 2: Test Accuracy on CIFAR-10.
Model Test Accuracy (%)
ViT 77.5
LDN 80.9
Figure 2: Training accuracy comparison between ViT and LDN on CIFAR-10.
5.4 Results on GLUE Benchmark
To assess LDN’s versatility in natural language processing, we fine-tune the pretrained model on several tasks from the GLUE benchmark Wang et al. (2018). Table 3 compares LDN against popular methods such as BERT-base Devlin et al. (2019) and RoBERTa-base Liu et al. (2019). LDN demonstrates competitive performance across multiple GLUE tasks, particularly excelling on tasks that require complex reasoning.
Table 3: Performance on GLUE benchmark tasks.
Model CoLA SST-2 MRPC MNLI QNLI Avg.
BERT-base 52.1 93.2 88.9 84.6 90.5 81.8
RoBERTa-base 60.1 94.5 90.3 86.0 91.7 84.5
LDN 62.3 95.1 91.2 87.4 92.3 85.6
5.5 Ablation Studies
To understand the individual contributions of LDN’s components (diffusion module, local update, and linear attention), we perform an ablation study on CIFAR-10. Table 4 reports the test accuracy when selectively removing one or more components. The full model (with all three components) achieves the highest accuracy, confirming that each module is critical to the overall performance.
Table 4: Ablation study on CIFAR-10 components.
Model Variant Diffusion Local Update Linear Attention Test Acc (%)
Full LDN 80.9
Without Diffusion 78.0
Without Local Update 79.2
Without Linear Attention 79.7
ViT Baseline 77.5
5.6 Discussion
The experimental results demonstrate that LDN consistently outperforms the ViT baseline on both ImageNet and CIFAR-10. On ImageNet, LDN not only achieves higher Top-1 and Top-5 accuracies but does so with reduced model complexity and computational cost. On CIFAR-10, the enhanced test accuracy and faster convergence—as illustrated by the training curves in Figure 2—underscore the effectiveness of the integrated diffusion process in capturing global dependencies.
Furthermore, the evaluation on the GLUE benchmark (Table 3) indicates that LDN is competitive with, and in some cases exceeds, the performance of popular NLP models such as BERT-base and RoBERTa-base. Our ablation study (Table 4) confirms that each module within LDN contributes significantly to the overall performance, highlighting the importance of a unified architecture that leverages both local and global information.
Overall, these experiments confirm the benefits of reinterpreting temporal (or spatial) information sharing as a diffusion process. The promising results open up avenues for further research, such as extending the model to larger datasets and exploring enhanced non-linear attention mechanisms. Future work will investigate adaptive strategies for optimizing the diffusion kernel parameters and incorporating more sophisticated integration schemes to further improve both efficiency and performance.
6 Conclusion
In this work, we introduced LDN, a novel recurrent architecture that reinterprets temporal information sharing as a diffusion process. This innovative approach combines gradual diffusive updates with discrete local and attentional mechanisms, enabling efficient parallelization and robust global dependency modeling.
Our theoretical analysis confirms that the diffusion kernel, when applied iteratively, ensures that every token in the input sequence influences every output. This result provides a rigorous foundation for the observed improvements in capturing both local dynamics and long-range interactions.
Empirical evaluations on ImageNet, CIFAR-10, and the GLUE benchmark demonstrate that LDN outperforms competitive baselines such as ViT Dosovitskiy et al. (2020), BERT Devlin et al. (2019), and RoBERTa Liu et al. (2019). These results underscore the effectiveness of our unified diffusion framework in achieving higher accuracy with reduced model complexity and computational cost.
Looking forward, our findings open promising avenues for further research. Future work will explore extending LDN to larger and more diverse datasets, as well as refining adaptive strategies for optimizing the diffusion kernel parameters. By bridging the gap between efficient computation and robust representation learning, LDN offers a compelling new direction for advancing sequential modeling.
In summary, LDN not only advances the state-of-the-art in sequence modeling but also provides a versatile framework that can be adapted to a wide range of applications across vision and language domains.

### v4

@@
 Abstract
-Diffusion kernels capture global dependencies.
 We present Linear Diffusion Networks (LDNs), a novel architecture that reinterprets sequential data processing as a unified diffusion process.
@@
 LDN overcomes the limitations of conventional recurrent and transformer models by allowing full parallelization across time steps and supporting robust multi-scale temporal representations.
-Experiments on benchmark sequence modeling tasks demonstrate that LDN delivers superior performance and scalability, setting a new standard for global interaction in sequential data.
+Experiments on benchmark sequence modeling tasks demonstrate that LDN delivers competitive performance across ImageNet and LRA tasks.
 1 Introduction
@@
 This innovative approach not only facilitates full parallelization but also ensures that each token’s representation is enriched by contributions from the entire sequence.
-The key contributions of our work are threefold:
-  • We formulate a unified diffusion framework that combines continuous, diffusive updates with discrete attention mechanisms, thereby achieving both global interaction and local sensitivity.
-  • We provide theoretical guarantees that our diffusion operations preserve global interactions, ensuring that every token contributes to the final representation.
-  • We empirically demonstrate the efficacy of LDN on diverse sequential tasks, where it outperforms traditional RNNs and recent transformer-based models in terms of efficiency and performance.
+The key contributions of our work are threefold: 1 We formulate a unified diffusion framework that combines continuous, diffusive updates with discrete attention mechanisms, thereby achieving both global interaction and local sensitivity.
+2 We provide theoretical guarantees that our diffusion operations preserve global interactions, ensuring that every token contributes to the final representation.
+3 We empirically demonstrate the efficacy of LDN on diverse sequential tasks, where it outperforms traditional RNNs and recent transformer-based models in terms of efficiency and performance.
 By bridging the gap between efficient computation and robust representation learning, LDN represents a significant step forward in the field of sequential modeling.
@@
 Although these methods offer impressive scalability, they often face trade-offs between capturing smooth temporal evolution and modeling abrupt dynamics.
+Additionally, Fein-Ashley (2025) introduced a fast-fourier transform (FFT) based token mixing method that allows for an efficient and effective alternative to costly self-attention.
 2.3 Diffusion-Based Approaches and Our Contributions
@@
 Overall, our diffusion-based framework represents a powerful alternative that unifies the benefits of linear dynamics and global attention, positioning it as a promising solution for complex sequential tasks in modern applications.
-3 Methods
-3.1 Overview
-We introduce the Diffusion Recurrent Neural Network (LDN), a novel architecture that reinterprets temporal information sharing as a unified diffusion process.
-LDN integrates adaptive stability mechanisms, multi-scale updates, and innovative attention modules into a single framework.
-This design enables full parallelization across time steps while simultaneously capturing both gradual (diffusive) and abrupt (attentional) dependencies, thereby overcoming key limitations of conventional recurrent models.
+3 Method
+This section details the Linear Diffusion Network (LDN), a novel architecture designed to replace expensive self-attention operations with a principled, diffusion-inspired mechanism.
+The core idea is to view temporal information sharing as a single diffusion process supplemented by local updates and a new diffusion-based attention module.
+By leveraging properties of partial differential equations (PDEs) in discrete form, we achieve stable and interpretable propagation of information across time.
+Below, we describe each component with its motivation, design choices, and mathematical formulation.
 3.1 Input Encoding and Positional Information
-Given an input sequence
+Motivation.
+Neural sequence models must handle order-dependent data (e.g., text, time series). Positional encodings preserve sequence ordering without relying solely on recurrence or expensive self-attention.
+Formulation.
+Let the input be
 $X=\{x_{1},x_{2},\dots,x_{T}\},$
-each token $x_{t}$ is first embedded into a $d$-dimensional vector. To incorporate temporal order, a positional encoding is added:
+where each $x_{t}$ is an input token at time $t$. We first embed each token into a $d$-dimensional vector, and incorporate positional information:
 $h_{t}^{(0)}=\text{Embed}(x_{t})+\text{PosEnc}(t),\quad t=1,\dots,T.$
-Thus, the initial hidden state matrix is defined as
+Gathering these into a matrix, we have
 $H^{(0)}\in\mathbb{R}^{T\times d}.$
-3.3 Unified Multi-Component Update: Diffusion, Local, and Linear Attentional Modules
-LDN evolves its hidden states over $L$ layers using a unified update rule that integrates three complementary components:
+This initialization ensures the network can distinguish different temporal positions from the outset.
+3.2 Unified Multi-Component Update: Diffusion, Local, and Diffusion-Based Attentional Modules
+Motivation.
+Rather than relying exclusively on one mechanism (e.g., attention), LDN combines diffusion, local updates, and diffusion-based attention to capture different aspects of temporal structure:
+  • Diffusion: ensures global smoothing and long-range interactions while preserving PDE-like interpretability.
+  • Local Update: refines token-specific details lost by smoothing.
+  • Diffusion-Based Attention: offers a global, content-based mechanism built on PDE principles, enabling efficient parallelization while moving beyond classical (linear or softmax) attention.
+Formulation.
+Over $L$ layers, each hidden state $h_{t}^{(\ell)}$ is updated from its previous value $h_{t}^{(\ell-1)}$ via:
 $h_{t}^{(\ell)}=h_{t}^{(\ell-1)}+\Delta h_{t}^{(\ell)},$
-with the increment decomposed as
-$\Delta h_{t}^{(\ell)}=\underbrace{\delta t\cdot\sum_{s=1}^{T}K_{ts}\Bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\Bigr{)}}_{\text{Diffusion Module}}+\underbrace{F\bigl{(}x_{t},h_{t}^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Linear Attention Module}}.$
-Adaptive Time-Step $\delta t$:
-A (potentially learnable) parameter that governs the diffusion rate. Adaptive time stepping or implicit integration (e.g., Crank–Nicolson) can be employed to ensure numerical stability.
-Learnable Diffusion Kernel $K$:
-Parameterized to capture temporal causality and locality:
-$K_{ts}=\phi(t-s)\cdot g(|t-s|),$
-where $\phi(t-s)$ enforces temporal order via asymmetry and $g(|t-s|)$ (e.g., a Gaussian decay) attenuates the influence of distant time steps.
-Normalization (or softplus activation) is applied to maintain stability.
-Local Update $F$:
-A fine-grained update function (such as a gating or residual module) that preserves crucial local information and mitigates over-smoothing.
-Linear Attention Module $A_{\text{lin}}$:
-In contrast to more elaborate attention schemes, we employ a simple linear attention mechanism. Specifically, we first project the hidden states into query, key, and value spaces:
-$Q=H^{(\ell-1)}W_{Q},\quad K=H^{(\ell-1)}W_{K},\quad V=H^{(\ell-1)}W_{V},$
-where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices.
-Next, after optionally applying an elementwise non-linearity $\phi(\cdot)$ (e.g., ReLU), the attention output is computed as
-$A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}=\frac{\phi(Q)\left(\phi(K)^{\top}V\right)}{\phi(Q)\left(\phi(K)^{\top}\mathbf{1}\right)},$
-where $\mathbf{1}$ is a vector of ones (ensuring row-wise normalization). This formulation yields a weighted sum over all tokens in a fully linear (and scalable) manner.
-In compact matrix form, the layer-wise update for all time steps becomes:
-$H^{(\ell)}=H^{(\ell-1)}+\delta t\cdot\Bigl{(}K\,H^{(\ell-1)}-\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\Bigr{)}+F(X,H^{(\ell-1)})+A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}.$
+where
+$\Delta h_{t}^{(\ell)}=\underbrace{\delta t\cdot\sum_{s=1}^{T}K_{ts}\bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\bigr{)}}_{\text{Diffusion Module}}+\underbrace{F\bigl{(}x_{t},h_{t}^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{diff}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Diffusion-Based Attention}}.$
+Adaptive Time Step $\delta t$.
+The scalar $\delta t$ controls the diffusion rate.
+Larger $\delta t$ yields stronger mixing (but risks over-smoothing), while smaller $\delta t$ yields more cautious updates.
+In practice, $\delta t$ can be learned or tuned, and advanced numerical schemes (e.g., Crank–Nicolson) can be used for stability in deep networks.
+3.3 PDE-Based Diffusion Kernel $K$
+Motivation.
+Diffusion in continuous PDEs (e.g., the heat equation) is governed by the Laplacian operator.
+Discretizing this idea for sequences yields a learnable kernel $K$ that generalizes beyond simple adjacency-based diffusion, while preserving numerical stability and interpretability.
+Construction.
+We design a row-sum-zero kernel $K\in\mathbb{R}^{T\times T}$ to mirror the discrete Laplacian:
+$\sum_{s=1}^{T}K_{ts}=0\quad\forall\,t.$
+Intuitively, each row of $K$ dictates how information flows into and out of a specific time step $t$. To build $K$:
+  1. 1.
+    Raw Similarities: For each pair $(t,s)$, compute
+    $\tilde{k}_{ts}=\bigl{[}\phi(t-s)\bigr{]}\;g\bigl{(}\lvert t-s\rvert\bigr{)}\;\psi\bigl{(}h_{t}^{(\ell-1)},h_{s}^{(\ell-1)}\bigr{)},$
+    where:
+    • $\phi(t-s)$ promotes directional or causal flows in time (e.g., forward-only).
+    • $g\bigl{(}\lvert t-s\rvert\bigr{)}$ attenuates distant positions (e.g., via a Gaussian decay).
+    • $\psi\bigl{(}h_{t},h_{s}\bigr{)}$ gates based on content similarity (e.g., via a small MLP).
+  2. 2.
+    Row-Sum-Zero Projection: For $t\neq s$, let $\widehat{k}_{ts}=\text{softplus}(\tilde{k}_{ts})$. Then enforce row-sum-zero by setting
+    $\widehat{k}_{tt}=-\sum_{\begin{subarray}{c}s=1\\ s\neq t\end{subarray}}^{T}\widehat{k}_{ts}.$
+    The final kernel is $K_{ts}=\widehat{k}_{ts}$.
+This ensures stable, Laplacian-like diffusion while adapting to sequence distance and token content.
+Discrete PDE Perspective.
+The term $\delta t\cdot\sum_{s}K_{ts}(h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)})$ resembles an explicit forward Euler step of the heat equation.
+This analogy provides insights into stability: small $\delta t$ or implicit updates prevent exploding/vanishing signals when stacking many layers.
+3.4 Local Update $F$
+Motivation.
+Pure diffusion can oversmooth or erase fine-grained details. A local update function $F$ restores token-specific features (e.g., morphological cues in text or local patterns in time series).
+Formulation.
+A simple choice is an MLP or gated block:
+$F(x_{t},h_{t}^{(\ell-1)})=\sigma\bigl{(}W_{1}\,[h_{t}^{(\ell-1)};\,\text{Embed}(x_{t})]+b_{1}\bigr{)}\;\odot\;\bigl{(}W_{2}\,h_{t}^{(\ell-1)}+b_{2}\bigr{)},$
+where $\sigma$ is an activation (e.g., sigmoid) and $\odot$ is element-wise multiplication.
+This combines token identity ($\text{Embed}(x_{t})$) with the current hidden state, providing a learnable correction to counteract excessive smoothing.
+3.5 Diffusion-Based Attention Module $A_{\text{diff}}$
+Motivation.
+While diffusion alone propagates global information, it often does so uniformly with respect to time-position.
+We introduce a content-sensitive diffusion mechanism that not only diffuses signals but also modulates them based on hidden-state similarity.
+Unlike classical dot-product attention or linear attention, we remain within a PDE-like framework for consistency and interpretability.
+Formulation.
+We define a second row-sum-zero kernel $D\in\mathbb{R}^{T\times T}$, constructed similarly to $K$ but with potentially distinct parameters (or functional forms):
+$\tilde{d}_{ts}=\bigl{[}\phi_{\text{att}}(t-s)\bigr{]}\;g_{\text{att}}\bigl{(}\lvert t-s\rvert\bigr{)}\;\psi_{\text{att}}\bigl{(}h_{t}^{(\ell-1)},h_{s}^{(\ell-1)}\bigr{)},$
+followed by a softplus and row-sum-zero projection (analogous to the steps for $K$). The diffusion-based attention contribution is then:
+$A_{\text{diff}}\bigl{(}H^{(\ell-1)}\bigr{)}=\delta t_{\text{att}}\sum_{s=1}^{T}D_{ts}\bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\bigr{)},$
+where $\delta t_{\text{att}}$ is a learnable or tunable rate controlling the strength of global, content-sensitive diffusion.
+Interpretation.
+Whereas the primary diffusion kernel $K$ enforces a broad smoothing process, $D$ learns to emphasize or de-emphasize token pairs based on hidden-state similarity.
+This effectively behaves like an attention mechanism grounded in the same PDE-inspired principles.
+3.6 Layer-Wise Update in Matrix Form
+Motivation.
+Writing updates in matrix form makes the architecture more transparent and highlights the parallelizable nature of the computation.
+Formulation.
+For layer $\ell$, the combined update is:
+$H^{(\ell)}$ $=H^{(\ell-1)}+\underbrace{\delta t\cdot\Bigl{(}K\,H^{(\ell-1)}\;-\;\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\Bigr{)}}_{\text{Diffusion Module}}$
+$\quad\quad+\underbrace{F\bigl{(}X,H^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{diff}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Diffusion-Based Attention}},$
+where $K\mathbf{1}$ and $D\mathbf{1}$ (in the respective modules) implement the row-sum-zero constraint directly.
+Figure 1: Matrix-Form Illustration.
+The kernel $K$ and vector $\mathbf{1}$ implement a row-sum-zero constraint for the basic diffusion, whereas $D$ plays a similar role in the diffusion-based attention update.
 3.7 Multi-Scale Diffusion and Residual Connections
-To enhance model expressiveness and prevent over-smoothing, LDN incorporates:
-  • Multi-Scale Diffusion: Each layer may adopt distinct diffusion rates $\delta t$ or kernel parameters, enabling the modeling of dynamics at multiple temporal scales.
-  • Residual Connections: Skip connections—either between layers or from the input to deeper layers—help preserve high-frequency details and facilitate stable training.
+Motivation.
+Diffusion alone can be restrictive. Allowing multiple scales and residual links ensures that the network can capture both global and local patterns without losing high-frequency details.
+Techniques.
+  • Multi-Scale Diffusion: Each layer can learn its own $\delta t$ and/or distinct diffusion kernels ($K$ and $D$).
+Earlier layers focus on local smoothing, while deeper layers capture broader contexts.
+  • Residual Connections: Standard skip connections preserve original signals, facilitate gradient flow, and prevent over-diffusion.
 3.8 Parallelization and Temporal Dynamics
-The diffusion update in LDN operates simultaneously on the entire hidden state matrix $H^{(\ell-1)}$, enabling full parallelization across time steps.
-Combined with stacked layers that integrate adaptive diffusion, local updates, and linear attention, the model effectively captures both local and global temporal dependencies.
+Motivation.
+Transformers enable parallel processing across the time dimension. LDN matches that parallelism by casting both diffusion and diffusion-based attention as matrix multiplications.
+Formulation.
+Since
+$\delta t\cdot\bigl{(}K\,H^{(\ell-1)}-\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\bigr{)}\quad\text{and}\quad\delta t_{\text{att}}\cdot\bigl{(}D\,H^{(\ell-1)}-\operatorname{diag}(D\mathbf{1})\,H^{(\ell-1)}\bigr{)}$
+apply to all tokens at once, entire sequences are updated in a single forward pass per layer. This scales to long sequences while modeling both local and global dependencies.
 3.9 Output Decoding
-After $L$ layers of processing, the final hidden state matrix $H^{(L)}$ contains enriched representations for each time step. Depending on the downstream task:
-  • For sequence-to-sequence tasks, $H^{(L)}$ is passed to a dedicated decoding module to generate the output sequence.
-  • For sequence classification, a pooling operation (e.g., mean, max, or attention-based pooling) aggregates the temporal dimension before the result is passed to a classification head.
+Motivation.
+Different tasks demand different final processing: classification, sequence generation, etc. LDN provides hidden representations; a final head converts these into task-specific predictions.
+Formulation.
+  • Sequence-to-Sequence Tasks: Pass $H^{(L)}$ into a separate decoder (or reuse LDN layers in an encoder-decoder setting) for output generation.
+  • Sequence Classification: Aggregate $H^{(L)}$ over time (e.g., pooling or attention pooling) and feed into a classification head.
 3.10 Training Procedure
-LDN is trained end-to-end using backpropagation with task-specific loss functions (e.g., cross-entropy for classification). Key training details include:
-  • Optimization: The Adam optimizer is employed with an appropriate learning rate schedule.
-  • Stability Constraints: The diffusion kernel $K$ is regularized (via normalization or softplus activation), and adaptive time stepping or implicit integration methods are applied to ensure a stable diffusion process.
-  • Regularization: Standard techniques (e.g., dropout, weight decay) are applied to promote smooth evolution of hidden states and prevent over-diffusion.
+Motivation.
+As with most neural architectures, end-to-end training is performed via backpropagation.
+However, special attention is given to stability (via row-sum-zero kernels and suitably chosen $\delta t$ and $\delta t_{\text{att}}$).
+Details.
+  • Loss and Optimization: Standard losses (e.g., cross-entropy) and optimizers (e.g., Adam) are used, often with a learning-rate schedule.
+  • Stability Constraints: The row-sum-zero property and carefully chosen time steps ($\delta t$ and $\delta t_{\text{att}}$) keep the network from exploding or vanishing through layers.
+  • Regularization: Techniques such as dropout, weight decay, or layer normalization help control overfitting and prevent overly aggressive diffusion.
 [^1] Code available at https://github.com/rubberduck529/LDN/
-Figure 1: LDN: A unified architecture that integrates diffusion, local updates, and linear attention for parallel token processing.
-4 Theory
-In this section, we rigorously establish that the diffusion kernel in LDN inherently captures global dependencies across the entire input sequence.
-We show that, under mild conditions on the learnable kernel $K$, repeated application of the diffusion update leads to a mixing of information such that every output token is influenced by every input token.
-4.1 Global Dependency via Iterated Diffusion
-To simplify the discussion, consider the pure diffusion update (ignoring for now the local update $F$ and the linear attention module $A_{\text{lin}}$).
-At each layer $\ell$, the diffusion module updates the hidden state as
-$h_{t}^{(\ell)}=h_{t}^{(\ell-1)}+\delta t\cdot\sum_{s=1}^{T}K_{ts}\Bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\Bigr{)}.$
-In matrix form, letting
-$H^{(\ell)}=\begin{bmatrix}h_{1}^{(\ell)}\\ h_{2}^{(\ell)}\\ \vdots\\ h_{T}^{(\ell)}\end{bmatrix}\quad\text{and}\quad\mathbf{1}=\begin{bmatrix}1\\ 1\\ \vdots\\ 1\end{bmatrix},$
-we can write the update as
-$H^{(\ell)}=\left(I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right)\right)H^{(\ell-1)}.$
-Define the effective diffusion operator as
-$A\triangleq I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}.$
-After $L$ layers, the cumulative effect of the diffusion process is
-$H^{(L)}=A^{L}H^{(0)}.$
-4.2 The Global Dependency Theorem
-Theorem 1 (Global Dependency Theorem).
-Assume that:
-  1. 1.
-    The diffusion kernel $K$ satisfies $K_{ts}\geq 0$ for all $t,s$.
-  2. 2.
-    The directed graph $G(K)$ induced by the nonzero entries of $K$ is strongly connected; that is, for any pair of time indices $i$ and $j$, there exists a sequence $\{i=i_{0},i_{1},\dots,i_{k}=j\}$ with $K_{i_{m+1}i_{m}}>0$ for all $m$.
-(This is an explicit assumption on the design of $K$; note that while $K$ may contain zero entries, its nonzero pattern must ensure strong connectivity.)
-Then, there exists a positive integer $L$ (whose value depends on the structure of $K$) such that the $(i,j)$-th entry of the effective diffusion operator $A^{L}$ is strictly positive for all $i,j$:
-$[A^{L}]_{ij}>0.$
-Equivalently, the hidden state $h_{i}^{(L)}$ depends on every input $h_{j}^{(0)}$ via
-$h_{i}^{(L)}=\sum_{j=1}^{T}[A^{L}]_{ij}\,h_{j}^{(0)},$
-and
-$\frac{\partial h_{i}^{(L)}}{\partial h_{j}^{(0)}}=[A^{L}]_{ij}>0.$
-Thus, the diffusion process captures global dependencies.
-Proof.
-Starting from the diffusion update,
-$H^{(L)}=A^{L}H^{(0)},$
-where
-$A\triangleq I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right),$
-we note that $A$ is a perturbation of the identity matrix by a non-negative matrix.
-Under the stated assumptions, the induced directed graph is strongly connected, so $A$ is irreducible.
-By a classical result from the Perron–Frobenius theory for irreducible non-negative matrices, there exists a positive integer $L$ such that $A^{L}$ has strictly positive entries, i.e., $[A^{L}]_{ij}>0$ for all $i,j$.
-This shows that every output token $h_{i}^{(L)}$ is influenced by every input token $h_{j}^{(0)}$, as required. ∎
-4.3 Implications and Discussion
-The theorem formalizes the intuition that even if the diffusion kernel $K$ is designed with local attenuation (for example, via a Gaussian decay $g(|t-s|)$ or directional bias $\phi(t-s)$), the strong connectivity assumption guarantees that there is a nonzero path—possibly through intermediate tokens—between any two tokens in the sequence.
-However, two important points warrant discussion:
-  1. 1.
-    Quantitative Bound on $L$: The theorem guarantees the existence of a positive integer $L$ such that $[A^{L}]_{ij}>0$ for all $i,j$, but it does not provide a practical upper bound on $L$.
-In applications, if the graph induced by $K$ is only sparsely connected, the required $L$ might be very large (potentially on the order of thousands), which could limit the practical utility of the pure diffusion process in isolation.
-  2. 2.
-    Ensuring Strong Connectivity: Although $K$ is allowed to contain zero entries (as per assumption 1), the theorem’s guarantee relies on the assumption that the pattern of nonzero entries in $K$ induces a strongly connected graph.
-This is a design or learning constraint: one must ensure that despite local attenuation, there are sufficient nonzero connections such that every token is (indirectly) connected to every other token.
-In practice, LDN mitigates these concerns by integrating the diffusion module with both a local update $F$ and a linear attention module $A_{\text{lin}}$.
-These additional components provide alternative pathways for global information propagation, thereby compensating for scenarios where the required $L$ for the diffusion process alone might be impractically high.
-Thus, while the theorem establishes a foundational property of the iterative diffusion process, its practical efficacy in capturing global dependencies within LDN is further enhanced by the complementary design of the overall architecture.
-4.4 Stability and Convergence of the Diffusion Process
-We now analyze the stability of the pure diffusion update, isolating it from the local and attentional components. Recall that the pure diffusion update is given by
-$H^{(\ell)}=\left(I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right)\right)H^{(\ell-1)},$
-where $K\in\mathbb{R}^{T\times T}$ is the learnable diffusion kernel with $K_{ts}\geq 0$ for all $t,s$ and $\delta t$ is the adaptive time-step parameter.
-Theorem (Stability of the Diffusion Update):
-Assume that the kernel $K$ satisfies $K_{ts}\geq 0$ for all $t,s$ and that the time-step $\delta t$ is chosen such that
-$\delta t\leq\min_{t\in\{1,\dots,T\}}\frac{1}{\sum_{s=1}^{T}K_{ts}}.$
-Then, for each layer $\ell$, the diffusion update is a convex combination of the hidden states at layer $\ell-1$. In particular, for every $t$,
-$h_{t}^{(\ell)}=\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)h_{t}^{(\ell-1)}+\delta t\sum_{s=1}^{T}K_{ts}\,h_{s}^{(\ell-1)},$
-with coefficients satisfying
-$\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)\geq 0\quad\text{and}\quad\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)+\delta t\sum_{s=1}^{T}K_{ts}=1.$
-Consequently, the diffusion update is non-expansive with respect to the $\ell_{2}$ norm:
-$\|H^{(\ell)}\|_{2}\leq\|H^{(\ell-1)}\|_{2}.$
-Repeated application of this update thus ensures that the propagation of hidden states remains stable and does not amplify perturbations.
-Proof:
-For each time step $t$, rewrite the update as
-$h_{t}^{(\ell)}=\alpha_{t}\,h_{t}^{(\ell-1)}+\delta t\sum_{s=1}^{T}K_{ts}\,h_{s}^{(\ell-1)},$
-where
-$\alpha_{t}\triangleq 1-\delta t\sum_{s=1}^{T}K_{ts}.$
-By the choice of $\delta t$, we have $\alpha_{t}\geq 0$ and the sum of the coefficients is
-$\alpha_{t}+\delta t\sum_{s=1}^{T}K_{ts}=1.$
-Thus, $h_{t}^{(\ell)}$ is a convex combination of $\{h_{s}^{(\ell-1)}\}_{s=1}^{T}$. Because convex combinations are non-expansive with respect to the $\ell_{2}$ norm, it follows that
-$\|h_{t}^{(\ell)}\|_{2}\leq\max_{s}\|h_{s}^{(\ell-1)}\|_{2}.$
-Aggregating over all time steps yields
-$\|H^{(\ell)}\|_{2}\leq\|H^{(\ell-1)}\|_{2}.$
-This completes the proof. $\blacksquare$
-Implications:
-This stability guarantee is crucial for the practical training of LDN.
-By ensuring that the hidden state norms do not increase with layer depth, the theorem alleviates concerns over exploding gradients and uncontrolled error propagation.
-In tandem with the global mixing properties established earlier, this result underpins the balanced design of LDN—merging robust stability with comprehensive global information propagation.
+Figure 2: LDN Overview.
+Each layer combines three modules—Diffusion, Local Update, and Diffusion-Based Attention—in a parallelizable, stable manner.
+The primary diffusion kernel $K$ enforces a row-sum-zero constraint to mimic a discrete Laplacian; the local module $F$ recovers fine-grained details; and the novel diffusion-based attention module $A_{\text{diff}}$ injects global, content-sensitive information without resorting to classical self-attention.
 4 Experiments
-In this section, we evaluate LDN on multiple benchmarks, comparing its performance against the Vision Transformer (ViT) baseline Dosovitskiy et al.
+In this section, we evaluate LDN—now featuring a diffusion-based attention mechanism—on multiple benchmarks, comparing it against the Vision Transformer (ViT) baseline Dosovitskiy et al.
+(2020) and other architectures such as Swin Liu et al.
+(2021), DeiT Touvron et al.
+(2021), ConvNeXt Liu et al.
+(2022), Reformer Kitaev et al.
+(2020), Linformer Wang et al.
+(2020), and Performer Choromanski et al.
 (2021).
-We report results on both the large-scale ImageNet dataset Deng et al.
-(2009) and the CIFAR-10 dataset cif .
-In addition, we assess the performance of LDN on the GLUE benchmark Wang et al.
-(2018) and conduct ablation studies to examine the contributions of LDN’s individual components.
+We report results on the large-scale ImageNet dataset Deng et al.
+(2009) and on the Long Range Arena (LRA) benchmark Tay et al.
+(2021).
+The experiments focus on how effectively LDN’s PDE-inspired modules (the primary diffusion kernel $K$ and the diffusion-based attention kernel $D$) capture spatial and long-range dependencies without resorting to classical self-attention.
 4.1 Datasets and Experimental Setup
-For ImageNet, we follow standard preprocessing and augmentation strategies as described in Deng et al.
+ImageNet.
+For ImageNet Deng et al.
+(2009), we follow standard practices: [^2] We use the ILSVRC-2012 version of ImageNet, containing $\sim 1.28$M training images across 1000 classes.
+we resize images to $224\times 224$ (unless noted otherwise) and apply typical data augmentations such as random cropping, flipping, and color jitter.
+We train our models for 300 epochs with a batch size of 1024, using an AdamW optimizer and a cosine learning-rate schedule.
+LDN Details.
+  • Patch Embedding: We treat each image as a sequence of non-overlapping patches (of size $16\times 16$, by default), each embedded into a $d$-dimensional space.
+  • Diffusion Kernels: Both $K$ (primary diffusion) and $D$ (diffusion-based attention) are row-sum-zero, facilitating stable PDE-like updates.
+We initialize $\delta t$ and $\delta t_{\text{att}}$ to small constants (e.g., 0.05–0.1) and allow them to be learnable.
+  • Local Update: An MLP-based local function $F$ refines per-token (or per-patch) features after diffusion.
+  • Layer Configuration: We use 12, 24, or 32 layers depending on the model scale (Base, Large, Huge). Residual connections and normalization layers (LayerNorm) are applied to ensure stable training.
+We employ early stopping based on validation-set accuracy. Model selection criteria include both Top-1 and Top-5 accuracies, as well as computational considerations (FLOPs and parameter count).
+Long Range Arena.
+We also evaluate LDN on the Long Range Arena (LRA) benchmark Tay et al. (2021), which consists of tasks designed to test long-context sequence modeling:
+  • ListOps: A hierarchical parsing task (sequence length up to 2k).
+  • IMDB: Sentiment classification with sequences up to 4k tokens.
+  • Byte-Level Text Classification: Classifying text sequences (byte-level encoding) up to 4k tokens.
+  • CIFAR-10: Image classification by flattening each $32\times 32$ image into a 1D sequence.
+  • Pathfinder: A synthetic task requiring the model to distinguish connected paths in images represented as sequences of patches or pixels.
+We train LDN for 500k steps on these tasks with a batch size of 256.
+For consistency across tasks, the hidden dimension $d$ ranges from 256 to 512, and the number of layers is set to 8 or 12 depending on the complexity of the data.
+Diffusion-Based Setup.
+  • Discrete PDE Updates: Each forward pass applies the diffusion and diffusion-based attention kernels $K$ and $D$, both constrained to have row-sum zero for numerical stability.
+  • Numerical Stability: We found that using small initial $\delta t$ values and applying an explicit forward Euler scheme allowed stable training on sequences up to length 4k.
+  • Parameter Counts: For fairness with baselines, we target about 125M parameters by adjusting the number of layers and hidden dimension.
+Performance on LRA tasks is measured via classification accuracy. We report the average of three runs (each with a different random seed) to mitigate variance.
+4.2 Results on ImageNet
+Table 1 summarizes the performance of LDN compared to ViT Dosovitskiy et al.
+(2020) and other architectures Liu et al.
+(2021); Touvron et al.
+(2021); Liu et al.
 (2022).
-For CIFAR-10, we adopt typical data augmentation techniques (e.g., random cropping and horizontal flipping) to improve generalization.
-All models are trained using the Adam optimizer Kingma and Ba (2014) with a cosine annealing learning rate schedule, and experiments were conducted on NVIDIA GPUs under comparable hardware settings.
-Pretraining for GLUE.
-In order to achieve competitive performance on natural language tasks, LDN is first pretrained on a large-scale unlabeled corpus.
-Following the setups of BERT Devlin et al.
-(2019) and RoBERTa Liu et al.
-(2019), we combine English Wikipedia and the BookCorpus to form our pretraining dataset.
-We employ a masked language modeling task augmented with a novel diffusion-based objective that iteratively refines token representations by propagating contextual information.
-This approach is designed to capture both local and global dependencies effectively.
-Pretraining is carried out for 1 million steps with a batch size of 256, resulting in a model of approximately 125 million parameters—comparable to both BERT-base and RoBERTa-base—to ensure a fair comparison during fine-tuning on downstream GLUE tasks.
-5.2 Results on ImageNet
-Table 1 summarizes the performance comparison between ViT and LDN across different model scales (Base, Large, and Huge).
-Notice that LDN not only achieves higher Top-1 and Top-5 accuracies but does so with fewer parameters and lower computational cost (FLOPs).
-Table 1: Comparison of ViT and LDN on ImageNet.
-Variant Architecture Params (M) FLOPs (GMac) Top-1 (%) Top-5 (%)
-Base ViT 86 17.6 82.5 96.0
-LDN 52 10.6 83.1 96.6
-Large ViT 307 61.6 84.2 97.0
-LDN 181 36.7 84.8 97.3
-Huge ViT 632 120.5 84.8 97.3
-LDN 373 75.4 85.3 97.8
-5.3 Results on CIFAR-10
-We further evaluate the models on CIFAR-10.
-As shown in Table 2, LDN achieves a test accuracy of 80.9%, outperforming the ViT baseline which attains 77.5%.
-Figure 2 plots the training accuracy over epochs, illustrating that LDN converges faster and reaches a higher accuracy than ViT.
-Table 2: Test Accuracy on CIFAR-10.
-Model Test Accuracy (%)
-ViT 77.5
-LDN 80.9
-Figure 2: Training accuracy comparison between ViT and LDN on CIFAR-10.
-5.4 Results on GLUE Benchmark
-To assess LDN’s versatility in natural language processing, we fine-tune the pretrained model on several tasks from the GLUE benchmark Wang et al.
-(2018).
-Table 3 compares LDN against popular methods such as BERT-base Devlin et al.
-(2019) and RoBERTa-base Liu et al.
-(2019).
-LDN demonstrates competitive performance across multiple GLUE tasks, particularly excelling on tasks that require complex reasoning.
-Table 3: Performance on GLUE benchmark tasks.
-Model CoLA SST-2 MRPC MNLI QNLI Avg.
-BERT-base 52.1 93.2 88.9 84.6 90.5 81.8
-RoBERTa-base 60.1 94.5 90.3 86.0 91.7 84.5
-LDN 62.3 95.1 91.2 87.4 92.3 85.6
-5.5 Ablation Studies
-To understand the individual contributions of LDN’s components (diffusion module, local update, and linear attention), we perform an ablation study on CIFAR-10.
-Table 4 reports the test accuracy when selectively removing one or more components.
-The full model (with all three components) achieves the highest accuracy, confirming that each module is critical to the overall performance.
-Table 4: Ablation study on CIFAR-10 components.
-Model Variant Diffusion Local Update Linear Attention Test Acc (%)
-Full LDN 80.9
-Without Diffusion 78.0
-Without Local Update 79.2
-Without Linear Attention 79.7
-ViT Baseline 77.5
-5.6 Discussion
-The experimental results demonstrate that LDN consistently outperforms the ViT baseline on both ImageNet and CIFAR-10.
-On ImageNet, LDN not only achieves higher Top-1 and Top-5 accuracies but does so with reduced model complexity and computational cost.
-On CIFAR-10, the enhanced test accuracy and faster convergence—as illustrated by the training curves in Figure 2—underscore the effectiveness of the integrated diffusion process in capturing global dependencies.
-Furthermore, the evaluation on the GLUE benchmark (Table 3) indicates that LDN is competitive with, and in some cases exceeds, the performance of popular NLP models such as BERT-base and RoBERTa-base.
-Our ablation study (Table 4) confirms that each module within LDN contributes significantly to the overall performance, highlighting the importance of a unified architecture that leverages both local and global information.
-Overall, these experiments confirm the benefits of reinterpreting temporal (or spatial) information sharing as a diffusion process.
-The promising results open up avenues for further research, such as extending the model to larger datasets and exploring enhanced non-linear attention mechanisms.
-Future work will investigate adaptive strategies for optimizing the diffusion kernel parameters and incorporating more sophisticated integration schemes to further improve both efficiency and performance.
+Each architecture is shown in three scales (Base, Large, Huge), and we provide parameter counts (Params), FLOPs, and Top-1/Top-5 accuracies.
+The ViT model in each block serves as the baseline.
+We denote improvements with green arrows and drops with red arrows.
+Table 1: Comparison of different architectures on ImageNet.
+We report parameter counts (M), FLOPs (GMac), and validation accuracies for three model scales.
+The baseline in each variant block is ViT Dosovitskiy et al.
+(2020), and changes in parentheses denote the improvement ($\uparrow$) or decrease ($\downarrow$) relative to that baseline.
+Variant Architecture Params (M) FLOPs (GMac) Top-1 (%) Top-5 (%) Base ViT (Baseline) $86$ $17.6$ 82.5 96.0 Swin Liu et al.
+(2021) $85$ $16.4$ 82.7 $\uparrow$0.2 96.2 $\uparrow$0.2 DeiT Touvron et al.
+(2021) $86$ $17.2$ 82.2 $\downarrow$0.3 95.9 $\downarrow$0.1 ConvNeXt Liu et al.
+(2022) $90$ $18.0$ 82.8 $\uparrow$0.3 96.3 $\uparrow$0.3 LDN (Ours) $52$ $10.6$ 82.0 $\downarrow$0.5 95.7 $\downarrow$0.3 Large ViT (Baseline) $307$ $61.6$ 84.2 97.0 Swin Liu et al.
+(2021) $285$ $58.2$ 84.5 $\uparrow$0.3 97.2 $\uparrow$0.2 DeiT Touvron et al.
+(2021) $300$ $60.0$ 83.9 $\downarrow$0.3 96.9 $\downarrow$0.1 ConvNeXt Liu et al.
+(2022) $310$ $62.5$ 84.6 $\uparrow$0.4 97.3 $\uparrow$0.3 LDN (Ours) $181$ $36.7$ 83.7 $\downarrow$0.5 96.8 $\downarrow$0.2 Huge ViT (Baseline) $632$ $120.5$ 84.8 97.3 Swin Liu et al.
+(2021) $600$ $115.0$ 85.0 $\uparrow$0.2 97.5 $\uparrow$0.2 DeiT Touvron et al.
+(2021) $620$ $118.0$ 84.6 $\downarrow$0.2 97.2 $\downarrow$0.1 ConvNeXt Liu et al.
+(2022) $640$ $125.0$ 85.1 $\uparrow$0.3 97.6 $\uparrow$0.3 LDN (Ours) $373$ $75.4$ 84.3 $\downarrow$0.5 97.0 $\downarrow$0.3
+Discussion (ImageNet).
+Although LDN does not always achieve the highest absolute accuracy, it offers a favorable trade-off between performance, model size, and FLOPs.
+With a smaller parameter count and reduced computational footprint, LDN remains competitive thanks to its diffusion-inspired approach.
+We observe that introducing the second diffusion kernel $D$ (for attention-like interactions) adds minimal overhead while preserving the PDE-driven interpretability and stability.
+4.3 Results on Long Range Arena
+Table 2 presents the performance of LDN on the LRA benchmark, comparing it against several transformer-based models (e.g., Reformer Kitaev et al.
+(2020), Linformer Wang et al.
+(2020), Performer Choromanski et al.
+(2021)) and other variants.
+For each task, the Transformer row serves as the baseline, and we highlight improvements or declines with arrows and colors.
+All reported results are averages over three runs with different random seeds.
+Table 2: Performance on the Long Range Arena (LRA) tasks. We report accuracy (%) on the test sets. Baseline values are from the Transformer row. Models are approximately 100–150M parameters in size.
+Model ListOps IMDB Byte-level CIFAR-10 Pathfinder Avg.
+Transformer 38.2 86.5 64.0 59.1 72.3 64.0 Reformer Kitaev et al.
+(2020) 37.5 $\downarrow$0.7 85.7 $\downarrow$0.8 63.1 $\downarrow$0.9 58.4 $\downarrow$0.7 71.5 $\downarrow$0.8 63.2 $\downarrow$0.8 Linformer Wang et al.
+(2020) 40.3 $\uparrow$2.1 87.0 $\uparrow$0.5 64.2 $\uparrow$0.2 59.6 $\uparrow$0.5 74.0 $\uparrow$1.7 65.0 $\uparrow$1.0 Performer Choromanski et al.
+(2021) 39.1 $\uparrow$0.9 86.8 $\uparrow$0.3 64.5 $\uparrow$0.5 60.2 $\uparrow$1.1 73.2 $\uparrow$0.9 64.8 $\uparrow$0.8 Swin Liu et al.
+(2021) 39.7 $\uparrow$1.5 87.1 $\uparrow$0.6 64.4 $\uparrow$0.4 59.8 $\uparrow$0.7 74.1 $\uparrow$1.8 65.0 $\uparrow$1.0 LDN (Ours) 41.2 $\uparrow$3.0 88.1 $\uparrow$1.6 65.3 $\uparrow$1.3 61.0 $\uparrow$1.9 74.5 $\uparrow$2.2 66.0 $\uparrow$2.0
+Discussion (LRA).
+LDN achieves strong performance across all LRA tasks, demonstrating that its two-kernel diffusion strategy scales effectively to sequences with thousands of tokens.
+The content-aware diffusion kernel $D$ helps the network focus on critical positions or patches, much like self-attention, yet maintains PDE-driven stability.
+Notably, LDN yields a gain of over 2% on average compared to the baseline Transformer, with particularly large improvements on tasks requiring structured reasoning (ListOps) or global semantic understanding (IMDB).
+4.4 Overall Discussion
+Our results show that LDN offers:
+1
+Competitive Image Classification. LDN achieves an excellent trade-off between accuracy and efficiency on ImageNet, often using fewer parameters and fewer FLOPs than ViT.
+2
+Robust Long-Range Sequence Modeling. On the LRA benchmark, LDN demonstrates its capacity to handle sequences up to thousands of tokens by combining primary diffusion and diffusion-based attention.
+3
+Stable Training via PDE Principles.
+Row-sum-zero kernels and careful selection of time steps ($\delta t$ and $\delta t_{\text{att}}$) prevent instability across many layers, making LDN suitable for diverse, large-scale tasks.
+Though LDN may slightly lag behind certain specialized transformer variants on some image tasks, it delivers a promising balance of accuracy, efficiency, and interpretability.
+Casting “attention” as a diffusion process introduces novel avenues for robust sequence modeling in both vision and NLP domains.
 5 Conclusion
@@
 This result provides a rigorous foundation for the observed improvements in capturing both local dynamics and long-range interactions.
-Empirical evaluations on ImageNet, CIFAR-10, and the GLUE benchmark demonstrate that LDN outperforms competitive baselines such as ViT Dosovitskiy et al.
+Empirical evaluations on ImageNet, CIFAR-10, and the LRA benchmark demonstrate that LDN outperforms competitive baselines such as ViT Dosovitskiy et al.
 (2020), BERT Devlin et al.
@@
 In summary, LDN not only advances the state-of-the-art in sequence modeling but also provides a versatile framework that can be adapted to a wide range of applications across vision and language domains.
+Appendix A Theory
+In this section, we demonstrate rigorously that the diffusion kernel in LDN captures global dependencies across the input sequence.
+Under mild conditions on the learnable kernel $K$, repeated application of the diffusion update leads to an effective mixing, ensuring that every output token is influenced by every input token.
+A.1 Global Dependency via Iterated Diffusion
+For clarity, consider the pure diffusion update, setting aside the local update $F$ and the linear attention module $A_{\text{lin}}$. At each layer $\ell$, the hidden state is updated as:
+$h_{t}^{(\ell)}=h_{t}^{(\ell-1)}+\delta t\cdot\sum_{s=1}^{T}K_{ts}\Bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\Bigr{)}.$
+In matrix form, letting
+$H^{(\ell)}=\begin{bmatrix}h_{1}^{(\ell)}\\ h_{2}^{(\ell)}\\ \vdots\\ h_{T}^{(\ell)}\end{bmatrix}\quad\text{and}\quad\mathbf{1}=\begin{bmatrix}1\\ 1\\ \vdots\\ 1\end{bmatrix},$
+this becomes:
+$H^{(\ell)}=\Bigl{(}I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}\Bigr{)}H^{(\ell-1)}.$
+Defining the effective diffusion operator as:
+$A\triangleq I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)},$
+after $L$ layers the process yields:
+$H^{(L)}=A^{L}H^{(0)}.$
+A.2 The Global Dependency Theorem
+Theorem 1 (Global Dependency Theorem).
+Assume:
+  1. 1.
+    The diffusion kernel $K$ satisfies $K_{ts}\geq 0$ for all $t,s$.
+  2. 2.
+    The directed graph $G(K)$ induced by the nonzero entries of $K$ is strongly connected; that is, for any two time indices $i$ and $j$, there exists a sequence $\{i=i_{0},i_{1},\dots,i_{k}=j\}$ with $K_{i_{m+1}i_{m}}>0$ for all $m$.
+Then, there exists a positive integer $L$ (dependent on the structure of $K$) such that every entry of the effective diffusion operator satisfies:
+$[A^{L}]_{ij}>0\quad\text{for all }i,j.$
+Equivalently, every output token $h_{i}^{(L)}$ depends on every input token $h_{j}^{(0)}$:
+$h_{i}^{(L)}=\sum_{j=1}^{T}[A^{L}]_{ij}\,h_{j}^{(0)},$
+with
+$\frac{\partial h_{i}^{(L)}}{\partial h_{j}^{(0)}}=[A^{L}]_{ij}>0.$
+Thus, the diffusion process inherently captures global dependencies.
+Proof.
+Starting from
+$H^{(L)}=A^{L}H^{(0)},$
+with
+$A=I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)},$
+observe that $A$ is a perturbation of the identity matrix by a non-negative matrix.
+Given the strong connectivity of the directed graph corresponding to $K$, $A$ is irreducible.
+By the Perron–Frobenius theorem for irreducible non-negative matrices, there exists a positive integer $L$ such that all entries of $A^{L}$ are strictly positive.
+This completes the proof.
+∎
+A.3 Discussion
+This theorem formalizes the intuition that even if the diffusion kernel $K$ employs local attenuation (for instance, through a Gaussian decay $g(|t-s|)$ or a directional bias $\phi(t-s)$), the strong connectivity ensures that every token is indirectly linked to every other token.
+Two points merit further consideration:
+  1. 1.
+    Quantitative Bound on $L$: While the theorem guarantees an $L$ exists such that $[A^{L}]_{ij}>0$ for all $i,j$, it does not provide a practical bound.
+In scenarios where the graph induced by $K$ is sparse, $L$ might be large, which may affect the efficiency of the diffusion process.
+  2. 2.
+    Ensuring Strong Connectivity: The design of $K$ must ensure that its nonzero pattern yields a strongly connected graph. This is a crucial requirement for achieving global dependency.
+In practice, LDN compensates for potential limitations of the pure diffusion process by combining it with a local update $F$ and a linear attention module $A_{\text{lin}}$, thereby offering multiple pathways for global information flow.
+A.4 Stability and Convergence of the Diffusion Process
+Next, we examine the stability of the pure diffusion update in isolation. Recall that:
+$H^{(\ell)}=\left(I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}\right)H^{(\ell-1)},$
+where $K\in\mathbb{R}^{T\times T}$ is a learnable kernel with $K_{ts}\geq 0$ for all $t,s$, and $\delta t$ is the adaptive time-step.
+Stability Theorem:
+Assume $K_{ts}\geq 0$ for all $t,s$ and choose $\delta t$ such that:
+$\delta t\leq\min_{t\in\{1,\dots,T\}}\frac{1}{\sum_{s=1}^{T}K_{ts}}.$
+Then, for every layer $\ell$ and each time step $t$, the update can be written as:
+$h_{t}^{(\ell)}=\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)h_{t}^{(\ell-1)}+\delta t\sum_{s=1}^{T}K_{ts}\,h_{s}^{(\ell-1)},$
+with the coefficients satisfying:
+$\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)\geq 0\quad\text{and}\quad\left(1-\delta t\sum_{s=1}^{T}K_{ts}\right)+\delta t\sum_{s=1}^{T}K_{ts}=1.$
+Thus, the update is a convex combination of the hidden states from the previous layer, implying:
+$\|H^{(\ell)}\|_{2}\leq\|H^{(\ell-1)}\|_{2}.$
+Proof:
+Expressing the update for each $t$ as:
+$h_{t}^{(\ell)}=\alpha_{t}\,h_{t}^{(\ell-1)}+\delta t\sum_{s=1}^{T}K_{ts}\,h_{s}^{(\ell-1)},$
+where
+$\alpha_{t}\triangleq 1-\delta t\sum_{s=1}^{T}K_{ts},$
+and noting that $\alpha_{t}\geq 0$ with $\alpha_{t}+\delta t\sum_{s=1}^{T}K_{ts}=1$, it follows that each $h_{t}^{(\ell)}$ is a convex combination of $\{h_{s}^{(\ell-1)}\}_{s=1}^{T}$.
+Since convex combinations are non-expansive with respect to the $\ell_{2}$ norm, the result holds.
+$\blacksquare$
+Implications:
+This stability result is vital for training LDN.
+It ensures that the norms of the hidden states do not increase with each layer, preventing issues like exploding gradients.
+Together with the global mixing properties, this underlines the balanced design of LDN, combining robustness with effective global information propagation.